AOV051

Hobart 12m

Disk VSN: HOB+0101/1600/1024| Data volume at beginning: 1029GB

  • 1812UT Experiment started late, first source 0433-422
  • 1824UT dbbc reconfigured
  • 1829UT 0154-244 missed
  • Autocorrelations have been funky, suspect file has been messed with? Ignored for now (Eloise/Warren)
  • 0325 - Schedule resumed after repairs to second stage amplifier (a power transistor had failed). No useful data before this time. (JMc)
  • No tsys readings are available - CAL controller ha failed (JMc).

Yarragadee 12m

Disk VSN: HOB+0095/1600| Data volume at beginning: 0GB

  • 1800UT Experiment started ok (Eloise/Warren)

~~DISCUSSION~~